Please help solve this equation:
Solve the system of equations using substitution. Be sure to write your answer as an ordered pair
y = 2x + 1
7x-y=26

Thanks!

Answer by stanbon(75887)

y = 2x + 1
7x-y=26
----
Substitute for "y" and solve for "x":
7x - (2x+1) = 26
5x = 27
x = 27/5
----
Since y = 2x+1
y = 2(27/5)+1
y = 54/5 +5/5
y = 59/5
----
Solution: (27/5 , 59/5)
===========================
Cheers,
Stan H.
================